OpenStudy (anonymous):

Which statement is true about the product square root of 2(5 + square root of 8)?

OpenStudy (anonymous):

@Purplerainbowcherry @RatzLover95 @Angel$18

OpenStudy (anonymous):

@sloths

OpenStudy (anonymous):

are there options?

OpenStudy (anonymous):

Yes, here they are It is rational and equal to 7. It is rational and equal to 9. It is irrational and equal to 4 + square root of 10. It is irrational and equal to 4 + 5square root of 2. Dont know why they didnt post.. haha

OpenStudy (anonymous):

I'm pretty sure its the last one

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

Need a little more info than, I think its the last one.. haha

OpenStudy (anonymous):

Can you help with a few more problems after this one?

OpenStudy (anonymous):

ok i'll try to explain best i can so you distribute the root 2 first which gives you\[\sqrt{10}+\sqrt{16}\] then you find the root of 16, which is 4 which then gives you \[4\sqrt{10}\] Then since 10 doesn't have a root you simplify it, giving you a final answer of \[4+5\sqrt{2}\]

OpenStudy (anonymous):

I'm a bit busy but i could take a look

OpenStudy (anonymous):

Gotcha. Sounds good.

OpenStudy (anonymous):

Simplify. 4 square root of 2 end root plus 7 square root of 2 end root minus 3 square root of 2 2 square root of 8 8 square root of 2 8 square root of 6 6 square root of 8

OpenStudy (anonymous):

Simplify. 4 square root of 2 end root plus 7 square root of 2 end root minus 3 square root of 2 2 square root of 8 8 square root of 2 8 square root of 6 6 square root of 8

OpenStudy (anonymous):

Those two

OpenStudy (anonymous):

they're the same question

OpenStudy (anonymous):

hold on ill look at them

OpenStudy (anonymous):

looks like this?\[4\sqrt{2}+7\sqrt{2}-3\sqrt{2}\]

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

sure. that one you just add the outside numbers basically. so its 4+7-3 which equals 8 and then you put the root 2 on the end. your answer is \[8\sqrt{2}\]

OpenStudy (anonymous):

Sweet! I put in the last problem I needed help with above. :) :p

OpenStudy (anonymous):

I only see two questions, the ones i already answered

OpenStudy (anonymous):

One second

OpenStudy (anonymous):

Which statement is true about the difference square root of 7 - square root of 28? It is rational and equal to -2. It is rational and equal to 0. It is irrational and equal to -2square root of 7. It is irrational and equal tosquare root of 7

OpenStudy (anonymous):

looks like this? \[\sqrt{7}-\sqrt{28}\]

OpenStudy (anonymous):

No,